# Blockchains for Internet of Things: Fundamentals, Applications, and Challenges

### v4

@@ II-3 Permissioned / Consortium Blockchains
Security problems in permissioned blockchains involve assessing the robustness of the consensus mechanism and the composition of the validator nodes. In these networks, identity management and authentication are critical aspects of security, as ensuring the legitimacy of participants is fundamental to the system’s integrity. Therefore, the implementation of robust access control mechanisms, including user authentication and role-based access control (RBAC), is indispensable. Moreover, smart contract security, data privacy, network security, and compliance with regulatory standards are important considerations in permissioned blockchains, particularly in industries like finance and supply chain management.
@@ III-3 Permissioned Blockchain with the Internet of Medical Things
Workflow Overview. In this example, Institutional Review Boards (IRBs) approve clinical trials; investigators oversee them, while researchers monitor subjects and evaluate changes. HFB refers to the Hyperledger Fabric Permissioned Blockchain, and all its members (owner of blockchain nodes) are authorized participants. Patients first register on the blockchain network (Step 9). They grant permissions using web interfaces (Step 8). Their sensitive data is stored in an AWS cloud database or AWS S3 bucket, with only authorized researchers able to access it (Step 10). Specifically, third-party users need to send requests to ask for permission from patients. After obtaining patient consent, the smart contract provides a data endpoint link to the researchers. A distributed threshold signature [14] facilitates clinical approval. When an IRB member creates a trial (Step 2), its details are broadcasted to all investigators (Step 1). After verification, investigators sign and move to a combination function (Step 3). If validated, the trial gets committed to the blockchain (Step 4). Researchers require permission to access patient trials (Steps 5, 6). Once they have the necessary approvals, they can download patient data for study. The complete implementation and evaluations are detailed in [15].

### v5

@@ II-3 Permissioned / Consortium Blockchains
Security problems in permissioned blockchains involve assessing the robustness of the consensus mechanism and the composition of the validator nodes. In these networks, identity management and authentication are critical aspects of security, as ensuring the legitimacy of participants is fundamental to the system’s integrity. Therefore, the implementation of robust access control mechanisms, including user authentication and role-based access control (RBAC) [12], is indispensable. Moreover, smart contract security, data privacy, network security, and compliance with regulatory standards are important considerations in permissioned blockchains, particularly in industries like finance and supply chain management.
@@ III-3 Permissioned Blockchain with the Internet of Medical Things
Workflow Overview. In this example, Institutional Review Boards (IRBs) approve clinical trials; investigators oversee them, while researchers monitor subjects and evaluate changes. HFB refers to the Hyperledger Fabric Permissioned Blockchain, and all its members (owner of blockchain nodes) are authorized participants. Patients first register on the blockchain network. They grant consent using web interfaces (Step 8). Their sensitive data is stored and registered in an AWS cloud database or AWS S3 bucket (Step 6&7), with only authorized users able to access it (Step 9). Specifically, third-party users need to send requests to ask for permission from patients. After obtaining patient consent (Step 1&4), the smart contract provides a data endpoint link to the researchers (Step 5). Once they have the necessary approvals, they can download patient data for study. The complete implementation and evaluations are detailed in [15]. Another example, when a clinician wants to conduct a clinical trial, he/she needs to submit a request to IRB, and IRB members approve a trial (Step 2&3).
@@ V Conclusion
Acknowledgments
This work was supported in part by the U.S. National Science Foundation under Grants CNS-2312139 and CNS-2332834.